Omar collected twice as many dollar bills as fives and one fewer ten than fives. If Omar collected $58, how many tens, fives, and ones did he get?
Answer by ikleyn(52835)

.
Let F be the number of five-dollar bills,


Then the number of dollar bills is 2F

and the number of ten-dollar bills is (F-1).


The "money" equation is

    5F + 2F + 10*(F-1)  = 58,


Simplify and solve for F:

    17F = 68

      F = 68/17 = 4.


Answer.  4 five-dollar bills;  2*4 = 8 dollar-bills  and  4-1 = 3 ten-dollars bills.


Check.   4*5 + 8*1 + 3*10 = 58 dollars.   ! Correct !